# A Foundational Individual Mobility Prediction Model based on Open-Source Large Language Models

### v2

@@ Abstract
Individual mobility prediction plays a key role in urban transport, enabling personalized service recommendations and effective travel management. It is widely modeled by data-driven methods such as machine learning, deep learning, as well as classical econometric methods to capture key features of mobility patterns. However, such methods are hindered in promoting further transferability and generalization due to limited capacity to learn mobility patterns from different data sources, predict in out-of-distribution settings (a.k.a “zero-shot”). To address the challenge, this paper proposes the foundational MoBLLM model based on lightweight open-source large language models (LLMs) fine-tuned on mobility datasets. We conduct extensive experiments on six real-world mobility datasets to evaluate its performance and transferability for diverse prediction tasks across varying times (years), space (cities), and situations (e.g., disruptions). The MoBLLM model achieves best F1 and accuracy performance in all datasets Compared with the state-of-the-art deep learning models. It also shows better transferability and cost performance than commercial LLMs. Further experiment reveals its high adaptability to different situations under network change, policy intervention, special events, and incidents. These results demonstrate the foundational capability of the MoBLLM model in predicting individual mobility, which contributes to precise and robust provision of personalized information for transportation management.
@@ 1 Introduction
Individual mobility prediction studies human mobility patterns at an individual level to forecast future movements. It facilitates a broad spectrum of applications in transportation domain, such as personalized services of POI or route recommendations and incentive mobility management (Ma and Zhang, 2022). Data-driven methods are dominant in this field including conventional Markov-based machine learning models (Gambs et al., 2012; Mathew et al., 2012; Lu et al., 2013) and deep learning (DL) models such as recurrent neural network (RNN) or Transformer (Feng et al., 2018; Kong and Wu, 2018; Xue et al., 2021). However, they are still limited to specialized training on the mobility dataset from a single city or data source, leading to low transferability of adapting to different travel contexts in different cities. It would be valuable to develop a foundational individual mobility prediction model capable of addressing these challenges.
@@
Recently, Large Language Models (LLMs) have made significant progress, far surpassing previous state-of-the-art models such as BERT (Devlin et al., 2019) and T5 (Raffel et al., 2020). The LLMs become the new foundational model in Natural Language Processing (NLP) domain. In Artificial Intelligence (AI) systems, a foundational model is any model that is trained on broad data (generally using self-supervision at scale) that can be adapted (e.g., fine-tuned) to a wide range of downstream tasks (Bommasani et al., 2022). As fine-tuning techniques advance, LLMs can be applied to developing foundational models for other various disciplines (Wang et al., 2025; Awais et al., 2025). Inspired by the concept of foundational models in AI, we define the foundational mobility prediction model as one trained on extensive mobility data, capable of adapting to diverse prediction tasks across varying time scales (years), spatial scales (cities), and situational conditions (such as disruptions and interventions).
@@
LLMs have also inspired a range of applications in urban mobility (Ma et al., 2025; Wang et al., 2024), particularly in mobility prediction tasks. Specifically, typical LLM based methods like In-Context Learning (ICL) (Brown et al., 2020) and Chain-of-Thought (CoT) (Wang et al., 2023a; Wei et al., 2022) enable modeling mobility prediction problems with natural language (Wang et al., 2023b; Liang et al., 2024; Feng et al., 2024a). However, hallucination issues (Huang et al., 2025) may arise when applying these methods, as their use of commercial LLMs lacks specialized training on mobility datasets. Therefore, the existing prompt engineering based methods are limited in generalizing prediction under different contexts (lack foundational capability). Besides, fine-tuning based on pretrained LLMs are applied to developing the individual mobility prediction framework (Li et al., 2024; Gong et al., 2024). They begin by aligning visited locations with specific intents, then leverage LLMs to predict the next location based on this enriched semantic context. From model validation perspectives, most of studies focus on prediction / accuracy but have not sufficiently demonstrated the LLMs’ capabilities of generalization and transferability across different data sources and prediction scenarios.
To fill the gap, we propose a foundational individual mobility prediction model MoBLLM based on open-source LLMs. It uses the pre-trained LLMs as the core module with generic prediction prompts and introduces instruction tuning (Zhang et al., 2023c) to enhance the LLMs’ capacities in adapting to different prompt writing styles of users and ensuring high transferability on different types of mobility data (e.g., check-in data). It adopts the Parameter Efficient Fine-Tuning (PEFT) method to realize the parameter and cost efficiency of training compared with the conventional full-parameter fine-tuning method. The key contributions include:
@@
    Propose a foundational, open-source LLM-based individual mobility prediction model, which can adapt to different data from multiple sources and cities, various individual mobility prediction tasks and even diverse prediction scenarios.
@@
    Introduce an efficient instruction generation approach to produce plentiful CoT based prompts for adapting to different writing styles of users.
@@
    Validate the proposed method through extensive experiments on different mobility datasets, including GPS location, AFC, and check-in data from different cities over the world. It compares against state-of-the-art DL and LLM models in aspects of prediction accuracy and transferability (space, time, and situation).
@@ 2 Related Work
The problem of individual mobility prediction consists of predicting an individual’s possible next location or trip information given their historical mobility sequence data (Ma and Zhang, 2022). In early stage, the Markov based methods are widely applied to modeling location transition probabilities, i.e, the next location prediction is based on several previous visited locations (Gambs et al., 2012; Mathew et al., 2012; Lu et al., 2013). Although these methods are intuitive and interpretable, they have low performance due to the difficulty in learning complex mobility patterns. Therefore, we focus on literature based on deep learning and LLM.
@@ 2.1 Deep Learning based Methods
Deep learning based methods commonly model individual mobility prediction by capturing the spatiotemporal features of an individual’s historical trip sequence. For example, Liu et al. (2016) models local temporal and spatial contexts by extending an recurrent neural network (RNN) layer with time-specific and distance-specific transition matrices. DeepMove (Feng et al., 2018) adopted a sequence to sequence (seq2seq)-based framework to capture the multi-level periodicity for mobility prediction from lengthy and sparse trajectories. The DeepTrip model (Zhang et al., 2023a) extended the DeepMove to predict the next trip information of the AFC metro data. Other RNN-based models adopted the Long Short-Term Memory (LSTM) architecture to model the prediction problem (Kong and Wu, 2018; Li et al., 2020; Chen et al., 2020).
@@ 2.2 Large Language Model based Methods
Common LLM-based methods for mobility prediction employ In-Context Learning (ICL) approaches to prompt LLMs with prediction task descriptions and historical mobility data, as well as Chain-of-Thought (CoT) to guide LLMs to capture complex semantic information. Such approaches enable LLM to make predictions in a zero-shot way without requiring additional training or gradient update. For example, LLM-Mob (Wang et al., 2023b) adopted ICL to realize zero-shot individual mobility prediction. Similarly, LLM-MPE (Liang et al., 2024) predicted human mobility under public events by leveraging the unprecedented ability of LLM to process textual data. AgentMove (Feng et al., 2024a) achieved a generalized prediction of mobility for any city worldwide by introducing POI data from cities to capture the knowledge of the world. LingoTrip Qin et al. (2025) prompted LLMs to capture trip purposes, choice preferences and choice variability from individual’s historical trip data to comprehensively infer the next trip location. While these methods exhibit promising application potential, the performance still has significant room for improvement since they use commercial LLMs trained on millions of general corpora without training on any mobility dataset. Meanwhile, the commercial LLMs lacking specialized knowledge may cause hallucination issues when answering domain-specific questions (Huang et al., 2025).
@@
Overall, the existing methods based on prompt engineering or fine-tuning show high potential in modeling the individual mobility prediction problem and yield a wide range of valuable applications. However, they are still insufficient to build a foundational model for individual mobility prediction due to the unsolved issues in adapting to diverse prediction tasks across varying spatiotemporal scales and conditional scenarios.
@@ 3.1 Problem Definition
Task 4 (next trip destination prediction): Given a user’s sequence of $n-1$ chronological trips: $\textbf{Tr}=(Tr_{1},Tr_{2},\dots,Tr_{n-1})$, the task is to predict the next destination station of $Tr_{n}$. A trip $Tr_{i}$ includes start time of trip $t_{o_{i}}$, day of the week $w_{i}$, origin station $s_{o_{i}}$, duration $dur_{i}$ of last stay occurring before $Tr_{i}$, and destination station $d_{i}$. We define end time of trip $t_{d_{i}}$, and then the stay duration is $dur_{i}=t_{o_{i}}-t_{d_{i-1}}$.
@@
Tasks 3 and 4 are generally defined in the studies conducted on AFC trip data. Here, we adopt an activity-based prediction framework assuming that a stay between two trips contains only one activity, as shown in Figures 1 and 1, which may benefit the model performance (Mo et al., 2021; Qin et al., 2025). Specifically, we define $d_{0}=home$ and $t_{0}=4:00AM$, which means the content of the first activity can be confirmed as staying home. The activity after the last trip of the day is not considered because it is exactly staying home the next day. The information of a trip mainly includes origin station $o$, destination station $d$, start time of trip $t_{o}$, end time of trip $t_{d}$, week of day $w$, which are commonly employed in trip prediction studies (Zhang et al., 2023b; Zhao et al., 2018). However, we extend a trip with the duration of the last stay defined in Task 4, supposing it is associated with certain activity information. Given the context of prediction on AFC trip data, we can align some notations of Task 3 with the trip definition where $t_{s_{i}}=t_{d_{i}}$, $t_{e_{i}}=t_{o_{i+1}}$, $l_{s_{i}}=d_{i}$ and $l_{e_{i}}=o_{i+1}$.
@@ 3.2 MoBLLM Framework
Figure 2: The overview of the proposed MoBLLM framework.
@@ 3.3 Prompt Template for Individual Mobility Prediction
First, the template briefly introduces the prediction tasks according to the definitions of base tasks in Section 3.1. To improve LLMs’ capability of understanding the context of mobility data, we provide detailed descriptions about the formatted data, which are adjusted due to the practical data usage of the corresponding tasks. The output format regulates the LLM’s output by asking it to provide its prediction and reason, facilitating quick processing of the text output. Such an approach of asking explanation not only improves LLMs’ prediction performance in mobility prediction but also increases the interpretability and reliability of the model (Wang et al., 2023b).
The thinking guidance is the key to carrying out the LLMs’ zero-shot learning in individual mobility prediction. It guides LLMs to capture mobility patterns from different time scales, i.e., the first and second aspects of thinking guidance shown in Figure 3. Then the third aspect reminds LLMs to learn the semantic information from the trips’ spatiotemporal context.

### v3

@@ Abstract
Individual mobility prediction plays a key role in urban transport, enabling personalized service recommendations and effective travel management. It is widely modeled by data-driven methods such as machine learning, deep learning, as well as classical econometric methods to capture key features of mobility patterns. However, such methods are hindered in promoting further transferability and robustness due to limited capacity to learn mobility patterns from different data sources, predict in out-of-distribution settings (a.k.a “zero-shot”). To address this challenge, this paper introduces MoBLLM, a foundational model for individual mobility prediction that aims to learn a shared and transferable representation of mobility behavior across heterogeneous data sources. Based on a lightweight open-source large language model (LLM), MoBLLM employs Parameter-Efficient Fine-Tuning (PEFT) techniques to create a cost-effective training pipeline, avoiding the need for large-scale GPU clusters while maintaining strong performance. We conduct extensive experiments on six real-world mobility datasets to evaluate its accuracy, robustness, and transferability across varying temporal scales (years), spatial contexts (cities), and situational conditions (e.g., disruptions and interventions). MoBLLM achieves the best F1 score and accuracy across all datasets compared with state-of-the-art deep learning models and shows better transferability and cost efficiency than commercial LLMs. Further experiments reveal its robustness under network changes, policy interventions, special events, and incidents. These results indicate that MoBLLM provides a generalizable modeling foundation for individual mobility behavior, enabling more reliable and adaptive personalized information services for transportation management.
@@ 1 Introduction
Individual mobility prediction studies human mobility patterns at an individual level to forecast future movements. It facilitates a broad spectrum of applications in transportation domain, such as personalized services of POI or route recommendations and incentive mobility management (Ma and Zhang, 2022). Data-driven methods are dominant in this field including conventional Markov-based machine learning models (2; W. Mathew, R. Raposo, and B. Martins (2012); X. Lu, E. Wetter, N. Bharti, A. J. Tatem, and L. Bengtsson (2013)) and deep learning (DL) models such as recurrent neural network (RNN) or Transformer (1; D. Kong and F. Wu (2018); H. Xue, F. Salim, Y. Ren, and N. Oliver (2021)). However, they are still limited to specialized training on the mobility dataset from a single city or data source, leading to low transferability of adapting to different travel contexts in different cities. It would be valuable to develop a foundational individual mobility prediction model capable of addressing these challenges.
@@
Recently, Large Language Models (LLMs) have made significant progress, far surpassing previous state-of-the-art models such as BERT (Devlin et al., 2019) and T5 (Raffel et al., 2020). The LLMs become the new foundational model in Natural Language Processing (NLP) domain. In Artificial Intelligence (AI) systems, a foundational model is any model that is trained on broad data (generally using self-supervision at scale) that can be adapted (e.g., fine-tuned) to a wide range of downstream tasks (Bommasani et al., 2022). As fine-tuning techniques advance, LLMs can be applied to developing foundational models for other various disciplines (Wang et al., 2025; Awais et al., 2025).
@@
LLMs have also inspired a range of applications in urban mobility (Ma et al., 2025; Wang et al., 2024b), particularly in mobility prediction tasks. Specifically, typical LLM based methods like In-Context Learning (ICL) (Brown et al., 2020) and Chain-of-Thought (CoT) (Wang et al., 2023a; Wei et al., 2022) enable modeling mobility prediction problems with natural language (Wang et al., 2023b; Liang et al., 2024; Feng et al., 2024a). However, hallucination issues may arise when applying these methods, as their use of commercial LLMs lacks specialized training on mobility datasets.
The hallucination is first defined in natural language generation as generated content that is either nonsensical or unfaithful to the provided source content (Huang et al., 2025), which can be similarly defined as unexpected LLMs’ outputs under mobility prediction context. Therefore, the existing prompt engineering based methods are limited in generalizing prediction under different contexts (lack foundational capability). Besides, fine-tuning based on pretrained LLMs are applied to developing the individual mobility prediction framework (Li et al., 2024; Gong et al., 2024). They begin by aligning visited locations with specific intents, then leverage LLMs to predict the next location based on this enriched semantic context. From model validation perspectives, most prior studies focus primarily on prediction / accuracy but have not sufficiently demonstrated the LLMs’ capabilities of transferability and robustness across different data sources and prediction scenarios.
Inspired by the notion of foundation models in artificial intelligence, we use the term foundational model in a transportation-science sense rather than in an architecture-centric AI sense. In this paper, a foundational mobility model refers to a reusable, general-purpose behavioral representation that supports multiple mobility-related prediction and policy-relevant tasks across heterogeneous cities, time periods, and contextual conditions.
Unlike traditional transportation models that are calibrated separately for each dataset, city, or policy scenario, a foundational mobility model aims to learn a shared representation of individual mobility behavior that can be transferred, adapted, and reused across contexts. The novelty lies not in proposing a new neural architecture, but in establishing a unified modeling substrate for mobility prediction that remains valid across spatial scales (cities), temporal scales (years), and situational conditions (e.g., disruptions, pricing, and interventions).
Under this definition, LLMs serve as a practical implementation mechanism for building such a foundation, because they provide a flexible representation of heterogeneous mobility data, traveler attributes, and contextual factors within a single transferable model. The key contributions include:
@@
    We introduce a general-purpose mobility foundation model for individual mobility prediction that provides a unified, transferable behavioral representation across heterogeneous data sources, cities, temporal periods, and situational contexts, enabling consistent modeling of mobility behavior beyond case-specific calibration.
@@
    We develop a scalable instruction-based adaptation framework that enables large language models to be efficiently specialized for diverse mobility prediction tasks and user-specific data formats, supporting flexible transfer across datasets without retraining from scratch.
@@
    We conduct extensive empirical evaluation on multi-city mobility datasets—including GPS trajectories, Automated Fare Collection (AFC) records, and check-in data—to systematically assess prediction accuracy, cross-city transferability, temporal generalization, and robustness to contextual variation, thereby establishing both the strengths and empirical boundaries of the proposed foundation model paradigm.
@@ 2 Related Work
The problem of individual mobility prediction consists of predicting an individual’s possible next location or trip information given their historical mobility sequence data (Ma and Zhang, 2022). In early stage, the Markov based methods are widely applied to modeling location transition probabilities, i.e, the next location prediction is based on several previous visited locations (2; W. Mathew, R. Raposo, and B. Martins (2012); X. Lu, E. Wetter, N. Bharti, A. J. Tatem, and L. Bengtsson (2013)). Although these methods are intuitive and interpretable, they have low performance due to the difficulty in learning complex mobility patterns. Therefore, we focus on literature based on deep learning and LLM.
@@ 2.1 Deep Learning based Methods
Deep learning based methods commonly model individual mobility prediction by capturing the spatiotemporal features of an individual’s historical trip sequence. For example, Liu et al. (2016) models local temporal and spatial contexts by extending an recurrent neural network (RNN) layer with time-specific and distance-specific transition matrices. DeepMove (1) adopted a sequence to sequence (seq2seq)-based framework to capture the multi-level periodicity for mobility prediction from lengthy and sparse trajectories. The DeepTrip model (Zhang et al., 2023b) extended the DeepMove to predict the next trip information of the AFC metro data. Other RNN-based models adopted the Long Short-Term Memory (LSTM) architecture to model the prediction problem (Kong and Wu, 2018; Li et al., 2020; Chen et al., 2020).
@@ 2.2 Large Language Model based Methods
Common LLM-based methods for mobility prediction employ In-Context Learning (ICL) approaches to prompt LLMs with prediction task descriptions and historical mobility data, as well as Chain-of-Thought (CoT) to guide LLMs to capture complex semantic information. Such approaches enable LLM to make predictions in a zero-shot way without requiring additional training or gradient update. For example, LLM-Mob (Wang et al., 2023b) adopted ICL to realize zero-shot individual mobility prediction. Similarly, LLM-MPE (Liang et al., 2024) predicted human mobility under public events by leveraging the unprecedented ability of LLM to process textual data. LingoTrip (Qin et al., 2025) prompted LLMs to capture trip purposes, choice preferences and choice variability from individual’s historical trip data to comprehensively infer the next trip location. Besides, agentic LLM frameworks integrating ICL, CoT and Retrieval-Augmented Generation (RAG) are applied to extending mobility prediction with more dimensions such as activity or intent inference. Specifically, AgentMove (Feng et al., 2024a) achieved a generalized prediction of mobility for any city worldwide by introducing POI data from cities to capture the knowledge of the world. ZHMF (Li et al., 2025) introduced a framework that integrates short-term trajectory records with structured long-term memory retrieved by RAG technology. LLMob (Wang et al., 2024a) adopted an agentic LLM framework that accounts for individual activity patterns and motivations. While these methods exhibit promising application potential, the performance still has significant room for improvement since they use commercial LLMs trained on millions of general corpora without training on any mobility dataset. Meanwhile, the commercial LLMs lacking specialized knowledge may cause hallucination issues when answering domain-specific questions (Huang et al., 2025).
@@
The reviewed mobility-related works based on LLMs are summarized in Table 1 according to the primary modeling methods, data and problem settings. Overall, the existing methods based on prompt engineering or fine-tuning show high potential in modeling the individual mobility prediction problem and yield a wide range of valuable applications. However, they are still insufficient to build a foundational model for individual mobility prediction due to the unsolved issues in adapting to diverse prediction tasks across varying spatiotemporal scales and conditional scenarios.
Table 1: Comparison between the LLM-based related Methods and MoBLLM.
Model Main Method Data Type Foundational Capability Task
LLM-Mob ICL and CoT Check-in, GPS trajectory Limited in accuracy Individual mobility prediction
LLM-MPE ICL and CoT Taxi ridership Limited in accuracy
Mobility prediction under public
events
LingoTrip ICL and CoT Metro trip Limited in accuracy Individual mobility prediction
LLMob Agentic LLM Check-in Limited in accuracy Trajectory generation
ZHMF Agentic LLM Check-in Limited in accuracy
Intent-aware, cold-start mobility
prediction
AgentMove Agentic LLM Check-in, GPS trajectory Limited in accuracy Individual mobility prediction
LIMP
Agentic LLM,
fine-tuning
Call detail records Limited in data scope Intent-aware mobility prediction
Mobility-LLM
ICL and CoT,
fine-tuning
Check-in Limited in data scope Intent-aware mobility prediction
MoBLLM
ICL and CoT,
fine-tuning
Check-in, GPS trajectory,
metro trip
Strong Individual mobility prediction
@@ 3.1 Problem Definition
Task 4 (next trip destination prediction): Given a user’s sequence of n1n-1 chronological trips: $\textbf{Tr}=(Tr_{1},Tr_{2},\dots,Tr_{n-1})$, the task is to predict the next destination station of $Tr_{n}$. A trip $Tr_{i}$ includes start time of trip $t_{o_{i}}$, day of the week $w_{i}$, origin station $s_{o_{i}}$, duration $dur_{i}$ of last stay occurring before $Tr_{i}$, and destination station $d_{i}$. We define end time of trip $t_{d_{i}}$, and then the stay duration is $dur_{i}=t_{o_{i}}-t_{d_{i-1}}$.
@@
Tasks 3 and 4 are generally defined in the studies conducted on AFC trip data. Here, we adopt an activity-based prediction framework assuming that a stay between two trips contains only one activity, as shown in Figures 1 and 1, which may benefit the model performance (Mo et al., 2021; Qin et al., 2025). Specifically, we define $d_{0}=home$ and $t_{0}=4:00AM$, which means the content of the first activity can be confirmed as staying home. The activity after the last trip of the day is not considered because it is exactly staying home the next day. The information of a trip mainly includes origin station $o$, destination station $d$, start time of trip $t_{o}$, end time of trip $t_{d}$, week of day $w$, which are commonly employed in trip prediction studies (Zhang et al., 2023a; Zhao et al., 2018). However, we extend a trip with the duration of the last stay defined in Task 4, supposing it is associated with certain activity information. Given the context of prediction on AFC trip data, we can align some notations of Task 3 with the trip definition where $t_{s_{i}}=t_{d_{i}}$, $t_{e_{i}}=t_{o_{i+1}}$, $l_{s_{i}}=d_{i}$ and $l_{e_{i}}=o_{i+1}$.
@@ 3.2 MoBLLM Framework
Figure 2: The overview of the proposed MoBLLM framework. Given 4 base task prompt templates, Step 1 generates multi-style semi-complete instructions without any mobility data by prompting advanced commercial LLMs such as GPT-4o mini. Step 2 constructs the instruction dataset by randomly sampling and assembling real-world mobility data with the semi-complete instructions. Step 3 obtains the MoBLLM model (for prediction) by a standard supervised fine-tuning pipeline that uses PEFT approach to train an open-source LLM on the instruction dataset.
Building upon the zero-shot prompting strategies of LLM-Mob (Wang et al., 2023b), which leveraged the general reasoning capabilities of commercial LLMs for mobility prediction, our work addresses a key limitation: these general-purpose models are not specialized for individual mobility, which constrains their predictive performance. While supervised fine-tuning methods such as PEFT provide a pathway to domain adaptation (Lu et al., 2025), they also risk catastrophic forgetting—where the model loses its ability to generalize to contexts beyond its fine-tuning dataset (Sun and Dredze, 2024). This raises concerns that fine-tuned LLMs may fail to transfer effectively to mobility data with distinct contexts (e.g., other cities or user populations) or structural types (e.g., different data formats or temporal resolutions).
To mitigate this risk and ensure robust transferability, we construct the instruction dataset with two deliberate strategies. First, we incorporate multi-type mobility data from multiple cities, enabling the LLM to capture structural patterns that are invariant across heterogeneous sources. Second, we normalize all location labels in each dataset to a universal integer index (0 to M1M-1, where $M$ is the number of unique locations). This abstraction intentionally removes dataset-specific semantics, compelling the model to focus on the underlying mechanisms of mobility dynamics—such as temporal regularities, sequential dependencies, and routine stability—rather than memorizing spatial identifiers. In doing so, MoBLLM leverages the compositional and sequential reasoning capacities of LLMs to internalize a generalized representation of human movement behavior. Although this abstraction may reduce sensitivity to spatial semantics, it substantially enhances the model’s ability to generalize across datasets with different spatial and contextual properties, which is the defining property of a foundational mobility model.
From a theoretical perspective, the generalization capability of MoBLLM can be interpreted through the lens of domain-invariant representation learning (Muandet et al., 2013; Arjovsky et al., 2019). By abstracting away spatial semantics through normalized indices, MoBLLM encourages the LLM to encode mobility sequences in a latent behavioral space that is less dependent on city-specific coordinate systems or naming conventions. In this space, prediction is guided primarily by learned transition dynamics and temporal regularities rather than explicit spatial memorization. This interpretation is conceptually aligned with ideas from causal and invariant representation learning (Schölkopf et al., 2021), in the sense that the model aims to capture behaviorally stable patterns—such as daily routines and temporal dependencies—that tend to persist across contexts. However, we emphasize that this remains a conceptual lens: no formal causal or invariance guarantees are claimed. Consequently, adaptation to unseen datasets is viewed as distributional alignment within this latent behavioral space, rather than proof of causal mechanism discovery.
@@ 3.3 Prompt Template for Individual Mobility Prediction
First, the template briefly introduces the prediction tasks according to the definitions of base tasks in Section 3.1. To improve LLMs’ capability of understanding the context of mobility data, we provide detailed descriptions about the formatted data, which are adjusted due to the practical data usage of the corresponding tasks. The output format regulates the LLM’s output by asking it to provide its prediction and reason, facilitating quick processing of the text output. Such an approach of asking explanation not only improves LLMs’ prediction performance in mobility prediction but also increases the interpretability and reliability of the model (Wang et al., 2023b). The thinking guidance is the key to carrying out the LLMs’ zero-shot learning in individual mobility prediction. It guides LLMs to capture mobility patterns from different time scales, i.e., the first and second aspects of thinking guidance shown in Figure 3. Then the third aspect reminds LLMs to learn the semantic information from the trips’ spatiotemporal context.